OpenStudy (anonymous):

Can you find the area of a hollow cylinder using the equation of a cylindrical shell?

OpenStudy (anonymous):

\[V = 2\pi(r)(thickness)(height)\]

OpenStudy (anonymous):

So, I am trying to figure out if the volume of a hollow cylinder is equivalent to the volume of a cylindrical shell used in integral calculus. Hallow Cylinder = \[(\pi)(height)((r _{o})^2-(r _{i})^2)\] Cylindrical Shell = \[2(\pi)(r _{i})(height)(thickness)\] The subscript "o" means outer-radius, and "i" means inter-radius

OpenStudy (irishboy123):

this doesn't really make any sense. to me at least. what are you trying to achieve?

OpenStudy (anonymous):

I think OP is trying to bridge whatever gap there appears to be between the shell and washer methods?

OpenStudy (anonymous):

Let's see if I'm understanding the question properly... Consider two cylinders with equal heights \(h\) and radii \(r\) and \(R\), where \(r<R\). Place the smaller inside the larger. You can compute the volume of the space between the two cylinders by subtracting to get the first formula you mentioned: \(V=\pi h(R^2-r^2)\).